# Learning Neural Contracting Dynamics: Extended Linearization and Global Guarantees

### v3

@@ 4.2 Results
Table 1 presents our results. ELCD performs the best in all but two datasets, where it performs second-best. These results shows the benefit of the increased expressiveness allowed by the skew symmetric component of our parametrization. EFlow does perform comparable to ELCD on the pendulum datasets, slightly beating us for the 8-link, 16-dimensional setting. As Eflow represents its vector field as the gradient of a function, it is arguably well-suited for the modeling of a mechanical system, which is inherently gradient like due to the underlying energy function.
The Rosenbrock dynamics are stiff and difficult to learn. In our training of SDD, we observed lack of stability and convergence, resulting in very large DTWDs. Hence, we report NaN for the results of the SDD models on the Rosenbrock dataset.
@@ 5 Limitations
Our method assumes that the underlying dynamics of the data trajectories are contracting. However, the diffeomorphic transform allows for a wide class of stable systems to be represented. Our method is currently implemented for trajectories that converge to the same fixed point. But, this can be overcome by manually changing the fixed point, depending on which trajectory the input data came from. Also, right now our method is limited to dynamics in n, but we imagine that an extension to more general manifolds should be possible.
@@ 6 Conclusions
In this paper, we introduce ELCD, the first neural network-based dynamical system with global contractivity guarantees in arbitrary metrics. The main theoretical tools are extended linearization, equilibrium contraction, and a converse contraction theorem. To allow for contraction with respect to more general metrics, we use a latent space representation with dimension of the latent space equal to the dimension of the data space and train diffeomorphisms between these spaces. We highlight key advantages of ELCD compared to NCDS as introduced in [5], including global contraction guarantees, expressible parameterization, and efficient inference. We demonstrate the performance of ELCD on high-dimensional LASA datasets and simulated Pendulum and Rosenbrock dynamics. Our method shows consistent performance across all datasets.
@@
This work was supported in part by the NSF Graduate Research Fellowship under grant 213931, the AFOSR under award FA9550- 22-1-0059, the NSF under grant no. 2229876, the Department of Homeland Security, and by IBM. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the NSF or its federal agency and industry partners.
@@ A.2 Comparisons to Other Models
where $\psi:{}^{d}\rightarrow{}^{d}$ is a diffeomorphism, $\Phi:{}^{d}\rightarrow\real$ is a convex potential function, and $G_{\psi}(x)=D\mkern-0.75mu{\psi}(x)^{\top}D\mkern-0.75mu{\psi}(x)$ is the induced Riemmanian metrc. The dynamics (18) then has the interpretation of being the steepest descent for $\Phi\circ\psi$ on the Riemmanian manifold defined by metric $G_{\psi}$. Specifically, in [29], the convex potential function is defined to be $\Phi(y)=\|y-y^{\star}\|_{2}$, where $y^{\star}=\psi(x^{\star})$ and the diffeomorphism is parametrized via a coupling layer [11]. Note that contraction properties of dynamics of these form were studied in [37]. Since $\Psi$ is chosen to be convex but not strongly convex, these dynamics are only weakly contracting, i.e., satisfy (2) with $c=0$.
@@
A.3 Additional Details on Datasets
In this section, we provide additional elaboration on the $n$-link pendulum dataset and the Rosenbrock dataset.
$n$-link pendulum: The $n$-link pendulum is a simple mechanical system which is the interconnection of $n$ rigid links under the force of gravity. We specifically assume that there is damping on each of these links, which makes the resulting dynamical system stable and almost all trajectories converge to the downright equilibrium point. In this case, the state of the pendulum is described by the $n$ pairs $(\theta,\dot{\theta}_{i})$, where $\theta$ is the angle of the $i$-th link and $\dot{\theta}_{i}$ is its angular velocity. Thus, this dynamical system is $2n$ dimensional.
@@ A.4 Model Details
A.5 Training details
All experiments are trained with a batch size of 100, for 100 epochs, with an Adam optimizer and learning rate of 1e-3. All computation is done on CPUs.

### v4

@@ 4.2 Results
Table 1 presents our results. ELCD performs the best in all tasks. These results shows the benefit of the increased expressiveness allowed by the skew symmetric component of our parametrization. These benefits are more apparent in the pendulum dataset. The oscillatory behavior of the pendulum dynamics is a product of complex eigenvalues in its Jacobian. Our model’s skew symmetric component is what enables it to learn the pendulum dynamics so well. A sample of demonstration and learned pendulum trajectories is shown in figure (4).
The Rosenbrock dynamics are stiff and difficult to learn. In our training of SDD, we observed lack of stability and convergence, resulting in very large DTWDs. Hence, we report NaN for the results of the SDD models on the Rosenbrock dataset. While our model performs the best, there is still room for improvement. Handling such dynamics with multiple time scales is still an open challenge.
Figure 4: Phase plots of the two link-pendulum trajectories (blue) and the trajectories produced by ELCD (red). Each column is a different trajectory. The top row is the first link and the bottom row is the second link. The x-axis is angle and the y-axis is angular velocity.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
@@ 5 Limitations
Our method assumes that the underlying dynamics of the data trajectories are contracting. However, the diffeomorphism allows for a wide class of stable systems to be represented. Our method is currently implemented for trajectories that converge to the same fixed point. But, this can be overcome by manually changing the fixed point, depending on which trajectory the input data came from. Also, right now our method is limited to dynamics in n, but we imagine that an extension to more general manifolds should be possible.
As we are motivated by applications in imitation learning and robotics, in this work, we are primarily interested in problems where the underlying systems should be robustly stable, especially away from training data. For this reason, we focus on learning dynamics which are guaranteed to be globally contracting and do not address learning other types of systems, such as those with multiple fixed points, limit cycles, or chaotic attractors. We imagine that extensions of ELCD could capture these richer dynamical behaviors by leveraging weaker notions of contraction including local contraction (i.e., contractivity in the region of attraction of a stable equilibrium), $k$-contraction (contraction of $k$-dimensional bodies) [43], transverse contraction [28] and contraction in the Hausdorff dimension [44]. Moreover, the notion of translation invariance mentioned could be studied using semicontraction theory [11], i.e., contraction to a subspace.
@@ 6 Conclusions
In this paper, we introduce ELCD, the first neural network-based dynamical system with global contractivity guarantees in arbitrary metrics. The main theoretical tools are extended linearization, equilibrium contraction, and a converse contraction theorem. To allow for contraction with respect to more general metrics, we use a latent space representation with dimension of the latent space equal to the dimension of the data space and train diffeomorphisms between these spaces. We highlight key advantages of ELCD compared to NCDS as introduced in [6], including global contraction guarantees, expressible parametrization, and efficient inference. We demonstrate the performance of ELCD on high-dimensional LASA datasets and simulated multi-link pendulum and Rosenbrock dynamics. Our method shows consistent performance across all datasets.
@@
This work was supported in part by the NSF Graduate Research Fellowship under grant 2139319, the AFOSR under award FA9550-22-1-0059, the NSF under grant no. 2229876, the Department of Homeland Security, and by IBM. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the NSF or its federal agency and industry partners.
@@ A.2 Comparisons to Other Models
where $\psi:{}^{d}\rightarrow{}^{d}$ is a diffeomorphism, $\Phi:{}^{d}\rightarrow\real$ is a convex potential function, and $G_{\psi}(x)=D\mkern-0.75mu{\psi}(x)^{\top}D\mkern-0.75mu{\psi}(x)$ is the induced Riemmanian metric. The dynamics (18) then has the interpretation of being the steepest descent for $\Phi\circ\psi$ on the Riemmanian manifold defined by metric $G_{\psi}$. Specifically, in [32], the convex potential function is defined to be $\Phi(y)=\|y-y^{\star}\|_{2}$, where $y^{\star}=\psi(x^{\star})$ and the diffeomorphism is parametrized via a coupling layer [12]. Note that contraction properties of dynamics of these form were studied in [42]. Since $\Psi$ is chosen to be convex but not strongly convex, these dynamics are only weakly contracting, i.e., satisfy (2) with $c=0$.
@@
A.3 Universal Representation
Our ELCD model equipped with a coupling layer-based transformation can represent any contracting dynamical system. It is known that any smooth nonlinear system with a hyperbolically stable fixed point can be exactly linearized inside its basin of attraction via a suitable diffeomorphism [25]. Effectively, our extended linearization parameterization is tasked with learning the stable linear part, while the coupling layers aim to learn and approximate this suitable diffeomorphism. As our ELCD parametrization can universally approximate any linear model, we simply need our coupling layer to universally approximate all diffeomorphisms. Indeed, [38] proved that the coupling layers are universal approximators for diffeomorphisms.
A.4 Additional Details on Datasets
In this section, we provide additional elaboration on the multi-link pendulum dataset and the Rosenbrock dataset.
Multi-link pendulum: The multi-link pendulum is a simple mechanical system which is the interconnection of $n$ rigid links under the force of gravity. We specifically assume that there is damping on each of these links, which makes the resulting dynamical system stable and almost all trajectories converge to the downright equilibrium point. In this case, the state of the pendulum is described by the $n$ pairs $(\theta_{i},\dot{\theta}_{i})$, where $\theta_{i}$ is the angle of the $i$-th link and $\dot{\theta}_{i}$ is its angular velocity. Thus, this dynamical system is $2n$ dimensional.
@@ A.5 Model Details
A.6 Training details
All experiments are trained with a batch size of 100, for 100 epochs, with an Adam optimizer and learning rate of $10^{-3}$. All computation is done on CPUs.
NCDS [6] trained the model and the encoder to output $x_{t+1}$ from $x_{t}$ using a single Euler integration step: $x_{t+1}=\phi^{-1}(\phi(x_{t})+dt*f(\phi(x_{t}))$. Their loss was the MSE between the predicted and true $x_{t+1}$. We empirically find better performance by training our model to directly predict the velocity vector $\dot{x}_{t+1}=\phi_{J}^{-1}(x_{t})f(\phi(x_{t}))$, where $\phi_{J}^{-1}(x_{t})$ is the Jacobian inverse of $\phi(x_{t})$. Calculating the gradient of $\phi_{J}^{-1}(x_{t})$ efficiently required a slight alteration to the original $\mathcal{M}$-flow [8] code, which we have included in our repository.